OpenStudy (anonymous):

the dimensions of a room are 16 x 14 x 10meters. there are four windows of 1.3 x 1 .4 m and two doors of 2m x 1m. what will be the cost of white washing the walls and painting the doors and windows.if the rate of white washing is rs5 per m^2 and the rate of painting is rs8 per m^2.

OpenStudy (anonymous):

first try to determine the area that needs to be washed. this would just be the area of all 4 walls minus the areas of the windows and doors. then use your rate equation for the rate of washing. likewise do the same for the area of the doors and use the rate equation for rate of painting.

OpenStudy (anonymous):

the Lateral Surface Area of walls= 2h (l + b) = 2 * 10 m (16 m + 14 m) = 600 m^2 area of 4 windows = 1.82 m^2 * 4 = 7.28 m^2 Area of 2 doors = 2 * 2m^2 = 4m^2 the area remaining to be painted = 600 - (7.28+4) = 588.72 m^2

OpenStudy (anonymous):

Cost of whitewashing wall = Rs.5 ( per m^2 ) Cost of whitewashing 588.72 m^2 of wall = 588.72 * Rs.5 = Rs.2943.6 Cost of painting wall = Rs.8 ( per m^2 ) Cost of painting 588.72 m^2 of wall = 588.72 * Rs.8 = Rs.4709.76